# OccamLLM: Fast and Exact Language Model Arithmetic in a Single Step

### v1

@@ 4.1 Simple Arithmetic Problems
The results of these evaluations are shown in Table 2. More detailed results, including relative error and results for 3 and 5 digit arithmetic, are shown in Appendix A. OccamLlama has $100.0\pm 0.0$% accuracy on all tasks, missing 0 out of 9000 problems. On the other hand, GPT 4o with Code Interpreter missed 3 out of the 1800 problems it faced (we tested on less problems to save cost) and had accuracy $99.8\pm 0.1$%.
Furthermore, GPT 4o with Code Interpreter generates on average more than 54 tokens to answer these problems, whereas our model uses OccamNet on the first forward pass. This means that, barring advanced decoding techniques such as speculative decoding [38], GPT 4o would need to be more than 50x faster than OccamLlama per forward pass to be comparable in overall generation speed on these tasks.
Figure 3: Accuracy of OccamLlama and baselines on reasoning tasks. Higher is better. OccamLlama achieves better performance than Llama 3 8B and GPT 3.5 Turbo on tasks with challenging arithmetic, reaches average accuracy above that of Llama 3 8B and GPT 3.5 Turbo and close to GPT 4o + Code.
@@
4.2 Mathematical Reasoning Problems
To test the performance of OccamLlama on longer generation tasks requiring a mix of reasoning and arithmetic capabilities, we evaluate our method and baselines on the following six benchmarks: AddSub [39], MultiArith [33], MATH401 [8], Single Eq [40], and SVAMP [41].
Because many of the arithmetic operations required in these datasets are relatively simple, we also create MultiArith Float, a modification of MultiArith in which we select problems which are arithmetically more challenging, while requiring similar levels of reasoning. To this end, we select prompts having input numbers that can be replaced with floats (e.g., $3.5$ feet or \$$\$39.95$ are reasonable but $3.5$ people is not). Furthermore, we sample input values from ranges larger than those appearing in the MultiArith dataset, in cases where it is reasonable. Float operations and larger additions and multiplications are more difficult for the baseline LLMs but do not make a difference for OccamLLM, so this dataset is particularly useful to show the advantages of this system.
Figure 3 shows the results of these evaluations. More detailed results are shown in Appendix A. On MultiArith Float and Math401, two datasets requiring challenging arithmetic, OccamLlama outperforms not only Llama 3 8B but also GPT 3.5 Turbo. On MultiArith Float, we find that OccamLlama matches the performance of GPT 4o. At the same time, most other datasets in this benchmark do not involve challenging arithmetic, meaning that Llama 3 8B is well suited to solve these tasks without assistance; most of the difficulty of these tasks lies in the reasoning rather than in the arithmetic computations. This is further supported by the fact that GPT 4o with Code Interpreter does not outperform, and in fact often underperforms, GPT 4o. As such, it is remarkable that OccamLlama can achieve comparable accuracy to Llama 3 8B even when it is trained on substantially different data and evaluated on tasks without challenging arithmetic.
@@ 4.2 Mathematical Reasoning Problems
Figure 4: Examples from Llama 3 8B Instruct and OccamLlama on arithmetic and a mathematical reasoning problem from the MultiArith Float dataset. In OccamLlama, the LLM performs reasoning, the switch predicts when to use OccamNet, and OccamNet performs arithmetic operations. Inputs to OccamNet are highlighted in purple. Outputs of OccamNet are highlighted in green.
@@ 4.3 Limitations
In most of our experiments, we use a single-layer OccamNet as the symbolic network, enabling evaluation of single-operation arithmetic problems. This sometimes poses a challenge on reasoning problems, when the base language model generates compound expressions containing more than one operation, such as $3+5+7=$ or $3\cdot 20\%$ where a single-layer OccamNet cannot evaluate. We attempt to overcome this by prompting Llama to break down compound expressions into multiple steps, but we find it difficult to coerce Llama to follow these instructions. Fortunately, we observed that these compound expressions were typically simple enough for the Llama to evaluate without OccamNet. As such, in our experiments, we trained OccamLLM switch to avoid using OccamNet for compound operations. In future work, we can explore other solutions such as integrating a two-layer OccamNet as the symbolic network. Another challenge is that Llama often generates expressions in fractions or percentages, and again, we found it difficult to coerce Llama to generate only decimals; fractions and percentages also constitute compound expressions and thus limits a single-layer OccamNet’s ability to perform helpful arithmetic. We found that these issues are particularly acute in the GSM8K [42] dataset, where the expressions generated by Llama are not prevalent in the switch training data, so sometimes incorrectly triggers OccamNet, thereby degrading performance. We discuss this phenomenon in more detail in Appendix A.
@@ 5 Discussion
We benchmarked our method on challenging arithmetic tasks, achieving 100% accuracy whereas GPT 4o achieves only 40% performance on average. We also benchmarked our method on mathematical reasoning tasks, demonstrating that the OccamLlama switch can accurately balance the LLM for reasoning and OccamNet for arithmetic, and outperforming Llama 3 8B Instruct on tasks involving challenging arithmetic. Our work could enable smaller LLMs to be as good as much much bigger LLMs in arithmetic; moreover, integrating OccamLLM with larger LLMs like GPT 4o could further improve their arithmetic abilities without relying on a code interpreter, but we leave this for future work. Furthermore, at present, OccamLLM may not integrate with more advanced decoding techniques such as speculative decoding [38, 43]. We hope to explore these avenues in future work.
@@ Appendix A Further Experiment Details and Results
For all training and evaluation runs, we used a single 32 GB NVIDIA Tesla V100 GPU. The LLM used for the OccamLLM system is Llama 3 8B Instruct. For each of the weight decoders and the switch described in 3.1.2 and 3.1.3, respectively, we used two-layer MLPs of input size 4096 (Llama-3-8B-Instruct hidden size), intermediate size 64 and final size equal to the number of weights in the corresponding OccamNet layer or switch.
@@ A.1 Training Dataset
We created a training dataset consisting of a 80,000 examples split in 40,000 single queries and 40,000 sequences of concatenated queries. In the first case, a single prompt of those described in 3.2 was sampled and formatted in the Llama 3 8B Instruct chat template, and the result of the expression was added as the output of the model. 40% of the sampled prompts correspond to simple arithmetic, concretely +,-,*,/. We sampled from various input value ranges. We sampled integers and floating point numbers uniformly from $(-20000,20000)$ and $(-1000,1000)$, integers only from $(-100,100)$ and $(-10,10)$ and floating points only from $(-1,1)$.
@@
Another 40% corresponds to complex arithmetic involving square roots, logarithms, exponentials, trigonometric functions and computing the power of one number to another one. For the square root and the logarithm, we sampled integers and floats uniformly from $(0.01,20000)$ and $(0.01,100)$. For the exponential, we sampled integers and floats from $(-10,10)$. For the powers, both base an exponents were sampled as integers or floats from $(1,25)$ and $(-6,6)$, respectively.
@@
The remaining 20% corresponds to reasoning prompts, either single or multi step problems. The inputs were sampled with various ranges, sometimes as floats and sometimes as integers, depending on the context of the problem. To train this model, the multiple-query examples never finished in a multi reasoning problem, as these would require more than one layer of OccamNet depth to be predicted accurately.
@@
We first iterated the 80,000 examples prepending "Answer = " to the assistant response, thus training OccamNet to predict the result after the "= ". Next, we validated the model on out-of-distribution examples, were it was forced to predict the result right at the beginning of the assistant response. We noticed that the accuracy on this task was improving during training, but after the full dataset was iterated it still didn’t perform as well as when evaluated in-distribution. Therefore, we continued to train the model using examples of the same dataset but with no appendix to the assistant response, and it rapidly learnt the new task. We stopped at 28,000 iterations in this second stage.
To train the switch, we create a dataset of 50,000 examples. Each example consists of a series of variable length input-output pairs sampled from a collection of prompts, which are formatted in the Llama 3 8B Instruct chat template in a conversational style. For each pair, the tokens previous to the numbers that should be computed using OccamNet, which are the ones that the switch should not route to the LLM, are marked with a 1, and all the rest are marked with a 0.
25% of these pairs are created by taking one of the arithmetic expressions described above as input, and using as output either “Answer = ”, the same expression repeated or nothing at all.
70% of the pairs come from a collection of manually created and labeled examples, which illustrate in which cases the switch should route to OccamNet and, importantly, in which cases it shouldn’t. This collection was deisgned to cover a wide variety of situations where the LLM might need to use OccamNet for computations. Furthermore, it includes cases where the LLM should avoid calling OccamNet because doing so would produce a bad prediction. This is the case, for example, of instances where the LLM attempts to add three numbers simultaneously. If it were to use the 1-layer OccamNet, which can take 2 inputs at most, the result would be incorrect.
The remaining 5% of the prompts come from multi-step reasoning problems. The output for these is not a full response, but only “The answer is ”.
@@
For the two-layer OccamNet run, we generated a large set of programatically generated prompts of the form 3+974=3+97$\cdot$-4=3 + 97 ⋅ - 4 =, with the Llama 3 8B chat template applied.
@@ A.2 Training Hyperparameters
For all 1-layer OccamNet training runs, we used a batch size of $1$, a learning rate of $46e-$ 4 and a weight decay parameter of $0.01$. We use the gradient accumulation technique to achieve an effective batch size of $8$. We used a constant learning rate scheduler. We use 1000 samples from OccamNet per token.
For the 2-layer OccamNet run, we use a learning $41e-$ 4, we used a batch size of $1$, a learning rate of $46.e-$ 4 and a weight decay parameter of $0.01$. We use the gradient accumulation technique to achieve an effective batch size of $8$. We used a constant learning rate scheduler. We use 50,000 samples from OccamNet per token.
@@ A.3 Prompting
For for the division arithmetic tasks, we found that the language models often did not return decimals. As such, we appended “Give the answer in decimals.” to these prompts. Similarly, for the trigonometric functions evaluations, we explicitly ask the language models to take the input as radians, by formatting the prompts as "cos(X rad) =".
@@
We did not use a system prompt, except for on the MATH401 dataset, where we noticed that GPT 4o was returning fractions instead of decimals. As such, on MATH401 we used the system prompt “Give your answer in decimals.”
@@
We experimented with OccamLlama prompts
@@ A.5 Experimental Results
Tables 4 and 5 show in more detail the accuracy of OccamLlama and other baselines on arithmetic and reasoning tasks. We measure accuracy as described in the main text.
We note here that on reasoning tasks with challenging arithmetic, such as Multiarith Float and MATH401, OccamLlama performs on-par with or even outperforms GPT 4o and GPT 4o Code. In fact, on MATH401, OccamLlama is nearly a 10 percentage points more accurate than Llama.
The only dataset on which OccamLlama struggles is GSM8K [42]. We beleive the reason for this is primarily because this data involves a large number of fractions and percentages, which Llama does not convert to decimal format. As such, Llama often calls OccamNet with expressions such as “multiplying by 3/4 gives” or “5% of this gives.” Because the switch is not trained on many examples like these in which the number is not in decimal format, it does not realize that OccamNet cannot be used in these cases. So, the switch triggers OccamNet, which is capable of performing the correct operation because it requires two steps. Future work could address this issue by training the router with more data on this type of situation or by training an OccamLlama with a two layer OccamNet.
Relative error is another important metric that complements accuracy. It measures by how much did the answer miss the true result. For two models with a similar accuracy metric, the relative error they achieve can be very different. Table 6 shows the relative error for the arithmetic experiments.
@@
Table 5: Accuracy on reasoning tasks. Higher is better. Bold indicates best non-GPT-4o performance.
@@
Average 85.7$\pm$0.5 53.3$\pm$0.7 82.1$\pm$0.6 86.5$\pm$0.5 95.2$\pm$0.3 93.4$\pm$0.7
Table 6: Relative error on arithmetic tasks. Higher is Better. Bold indicates best performance.
@@
Appendix B Alternative Architectures and Losses
B.1 Alternative Architectures

### v2

@@ 4.1 Simple Arithmetic Problems
The results of these evaluations are shown in Table 2. More detailed results, including relative error and results for 3 and 5 digit arithmetic, are shown in Appendix A. OccamLlama has $100.0\pm 0.0$% accuracy on all tasks, missing 0 out of 9000 problems. On the other hand, we tested GPT 4o with Code Interpreter on fewer problems to save cost, and it missed 3 out of the 1800 problems it faced, achieving an accuracy of $99.8\pm 0.1$%.
Furthermore, GPT 4o with Code Interpreter generates on average more than 54 tokens to answer these problems, whereas our model uses OccamNet on the first forward pass. This means that, barring advanced decoding techniques such as speculative decoding [38], GPT 4o would need to be more than 50x faster than OccamLlama per forward pass to be comparable in answer generation speed on these tasks.
Figure 3: Accuracy of OccamLlama and baselines on mathematical problem solving tasks. Higher is better. OccamLlama achieves accuracy comparable to Llama 3 8B on benchmarks with simple arithmetic, higher accuracy than GPT 4o and GPT 4o + Code on on tasks with challenging arithmetic, and accuracy above Llama 3 8B and similar to GPT 3.5 Turbo on average.
@@
4.2 Mathematical Problem Solving
To test the performance of OccamLlama on more general mathematical problem solving tasks, we evaluate our method and baselines on the following six benchmarks: AddSub [39], GSM8K [40], MultiArith [33], MATH401 [8], Single Eq [41], and SVAMP [42]. All but MATH401 are word problems requiring longer generation and a mix of reasoning and arithmetic capabilities. MATH401 also includes multistep arithmetic problems which require more than one call to OccamLlama.
Because many of the arithmetic operations required in these datasets are relatively simple, we also create MultiArith Float, a modification of MultiArith in which we select problems which are arithmetically more challenging, while requiring similar levels of reasoning. To this end, we select prompts having input numbers that can be replaced with floats. For instance, $3.5$ feet or \$$\$39.95$ are reasonable but $3.5$ people is not. Furthermore, we sample input values from ranges larger than those appearing in the MultiArith dataset, in cases where it is reasonable. Float operations and larger additions and multiplications are more difficult for the baseline LLMs but do not make a difference for OccamLLM, so this dataset is particularly useful to show the advantages of the system we propose.
Figure 3 shows the results of these evaluations. More detailed results are shown in Appendix A. On MultiArith Float and MATH401, two datasets requiring challenging arithmetic, OccamLlama outperforms not only Llama 3 8B but also GPT 4o and GPT 4o + Code. At the same time, most other datasets in this benchmark do not involve challenging arithmetic, meaning that Llama 3 8B is well suited to solve these tasks without assistance; most of the difficulty of these tasks lies in the reasoning rather than in the arithmetic computations. This is further supported by the fact that GPT 4o with Code Interpreter never substantially outperforms and sometimes underperforms GPT 4o. As such, it is remarkable that OccamLlama can achieve comparable accuracy to Llama 3 8B even when it is trained on very different data and evaluated on tasks without challenging arithmetic. The only dataset for which OccamLlama performs substantially worse than Llama 3 8B is GSM8K, but we attribute this to an imperfect OccamLlama switch, likely stemming from text which is outside of the switch training distribution (see Section 4.3).
@@ 4.2 Mathematical Problem Solving
Figure 4: Examples from Llama 3 8B Instruct and OccamLlama on (top) an arithmetic problem and (bottom) a mathematical reasoning problem from the MultiArith Float dataset. In OccamLlama, the LLM performs reasoning, the switch predicts when to use OccamNet, and OccamNet performs arithmetic operations. OccamNet’s inputs and outputs are highlighted in purple and green, respectively.
@@ 4.3 Limitations
In our experiments, we use a single-layer OccamNet as the symbolic network, enabling evaluation of single-operation arithmetic problems. This sometimes poses a challenge on reasoning problems, when the base language model generates compound expressions containing more than one operation, such as $3+5+7=$. A single-layer OccamNet cannot evaluate these expressions. We attempt to overcome this by prompting Llama to break down compound expressions into multiple steps, but we find it difficult to coerce Llama to follow these instructions. Another challenge is that Llama often generates expressions in fractions or percentages, which also constitute compound expressions that are not properly handled by the OccamLLM system. Fortunately, we observed that these compound expressions were typically simple enough for the LLM to evaluate without OccamNet. Therefore, in our experiments, we trained the OccamLLM switch to avoid using OccamNet for compound operations. In future work, we can explore other solutions such as integrating a two-layer OccamNet as the symbolic network. We found that these issues are particularly acute in the GSM8K [40] dataset, where the expressions generated by Llama are not prevalent in the switch training data, causing it to sometimes incorrectly trigger OccamNet and degrade performance.
@@
These techniques demonstrate a design paradigm of OccamLlama: by tuning the behaviors of OccamNet and the switch, we can often avoid finetuning the LLM.
@@ 5 Discussion
We benchmarked our method on challenging arithmetic tasks, achieving 100% accuracy while GPT 4o achieves only 40% performance on average. We also benchmarked our method on mathematical problem solving tasks, demonstrating that the OccamLlama switch can accurately balance the LLM for reasoning and OccamNet for arithmetic, outperforming even GPT 4o and GPT 4o with Code Interpreter on tasks involving challenging arithmetic. Our work could enable smaller LLMs to be as performant as much larger LLMs in arithmetic. Moreover, integrating OccamLLM with larger LLMs like GPT 4o could further improve their arithmetic abilities without requiring a code interpreter, but we leave this for future work. Furthermore, at present, OccamLLM may not integrate with more advanced decoding techniques such as speculative decoding [38, 43]. We hope to explore these avenues in future work.
@@ Acknowledgements
The authors acknowledge the MIT SuperCloud and Lincoln Laboratory Supercomputing Center for providing HPC resources that have contributed to the research results reported within this paper.
@@ Appendix A Further Experiment Details and Results
For all training and evaluation runs, we used a single 32 GB NVIDIA Tesla V100 GPU. The LLM used for the OccamLLM system is Llama 3 8B Instruct.
In the experiments presented in Section 4, for each of the weight decoders and the switch, we used two-layer MLPs of input size 4096 (Llama 3 8B Instruct hidden size), intermediate size 64 and final size equal to the number of weights in the corresponding OccamNet layer or switch.
In the two-layer experiments presented in Section 6, for each of the weight decoders, we used two-layer MLPs of input size 4096 (Llama 3 8B Instruct hidden size), intermediate size 512, and final size equal to the number of weights in the corresponding OccamNet layer. We did not train a switch for this experiment as we did not test long-form generations.
@@ A.1.1 OccamNet Decoder
To train the OccamNet decoder, we created a training dataset consisting of a 80,000 examples split in 40,000 single queries and 40,000 sequences of concatenated queries. In the first case, we sampled a single prompt of those described in 3.2 and formatted it using the Llama 3 8B Instruct chat template. In the second case, we concatenated multiple prompts described in 3.2 without the chat template.
40% of the sampled prompts correspond to simple arithmetic, concretely +,,×+,-$\times$+ , - , ×, and $\div$. We sampled from various input value ranges, chosen at random: integers in $[-10,10]$, integers in $[-100,100]$, integers in $[-1000,1000]$, integers in $[-20000,20000]$, floating numbers in $[-1,1]$, and floating point numbers in $[-1000,1000]$.
@@
Another 40% corresponds to complex arithmetic involving square roots, logarithms, exponentials, trigonometric functions and computing one number to the power of another. For the square root and the logarithm, we sampled integers uniformly in either $[1,100]$ or $[1,20000]$ and floats uniformly in either $[0.01,100]$ or $[0.01,20000]$. For the exponential, we sampled integers and floats in $[-10,10]$. For the powers, we sampled the base as either an integer in $[1,25]$ or a float in $[0.1,25]$ and the exponent as an integer in $[-6,6]$.
@@
The remaining 20% corresponds to single or multi step problems reasoning prompts. The inputs were sampled with various ranges, sometimes as floats and sometimes as integers, depending on the context of the problem. Because a single-OccamNet-layer OccamLlama cannot solve a multi-step reasoning problem in a single step, we never end the multiple-query examples with a multi-step reasoning problem.
@@
We first iterated the 80,000 examples, prepending “Answer = ” to the assistant response, thus training OccamNet to predict the result after the “=”. Next, we validated the model on out-of-distribution examples where “Answer = ” was not appended. We noticed that the accuracy on this task was improving during training, but after the full dataset was iterated it still didn’t perform as well as when evaluated in-distribution. Therefore, we continued to train the model using examples of the same dataset but with no “Answer = ” at the beginning of the assistant response. The model rapidly learned the new task. We stopped at 28,000 iterations of this second stage.
@@
For the two-layer OccamNet run, we generated a large set of programmatically generated prompts of the form 3+974=3+97$\cdot$-4=3 + 97 ⋅ - 4 =, with the Llama 3 8B chat template applied.
A.1.2 Switch Decoder
To train the switch decoder, we created a dataset of 50,000 examples. Each example consists of a series of variable length input-output pairs sampled from a collection of prompts, which are formatted in the Llama 3 8B Instruct chat template in a conversational style. For each pair, the tokens previous to the numbers that should be computed using OccamNet, which are the ones that the switch should not route to the LLM, are marked with a 1, and all the rest are marked with a 0.
25% of these pairs are created by taking one of the arithmetic expressions described above as input, and using as assistant output either “Answer = ”, the same expression repeated with an “=”, or nothing at all.
70% of the pairs come from a collection of manually created and labeled examples, which illustrate in which cases the switch should route to OccamNet and, importantly, in which cases it shouldn’t. This collection was designed to cover a wide variety of situations where the LLM might need to use OccamNet for computations. Furthermore, it includes cases where the LLM should avoid calling OccamNet because doing so would produce a bad prediction. This is the case, for example, of instances where the LLM attempts to add three numbers simultaneously. If it were to use the 1-layer OccamNet, which can take 2 inputs at most, the result would be incorrect.
The remaining 5% of the prompts come from multi-step reasoning problems. The output for these is not a full response, but only “The answer is ”. In such cases, a single-OccamNet-layer OccamLlama cannot solve the problem, so the switch output is set to 0.
@@ A.2 Training Hyperparameters
For all 1-layer OccamNet training runs, we used a batch size of $1$, a learning rate of $46e-$ 4 and a weight decay parameter of $0.01$. We use gradient accumulation to achieve an effective batch size of $8$. We used a constant learning rate scheduler. We take 1000 samples from OccamNet per token.
For the 2-layer OccamNet run, we used a batch size of $1$, a learning rate of $41e-$ 4 and a weight decay parameter of $0.01$. We use the gradient accumulation technique to achieve an effective batch size of $8$. We used a constant learning rate scheduler. We take 50,000 samples from OccamNet per token.
@@ A.3 Prompting
For the division arithmetic tasks, we found that the language models often did not return decimals. As such, we appended “Give the answer in decimals.” to these prompts. Similarly, for the trigonometric functions evaluations, we explicitly ask the language models to take the input as radians, by formatting the prompts as "cos(X rad) =".
@@
We did not use a system prompt, except for the MATH401 dataset, where we noticed that GPT 4o was returning fractions instead of decimals. As such, on MATH401 we used the system prompt “Give your answer in decimals.”
@@
We experimented with OccamLlama prompts, but discovered that not including a system prompt was most effective.
@@ A.5 Experimental Results
Tables 4 and 5 show in more detail the accuracy of OccamLlama and other baselines on arithmetic and mathematical problem solving tasks. We measure accuracy as described in the main text.
We note here that on datasets with challenging arithmetic, such as Multiarith Float and MATH401, OccamLlama outperforms GPT 4o and GPT 4o Code. In fact, on MultiArith Float, OccamLlama 20 percentage points more accurate than GPT 4o and and 40 percentage points more accurate than Llama 3 8B. Similarly, on MATH401, OccamLlama is nearly 10 percentage points more accurate than GPT 4o and 25 percentage points more accurate than Llama 3 8B. Although MATH401 does not include word problems, it does include some arithmetic expressions that require multiple calls to OccamNet to solve, meaning it requires both reasoning (to determine how to break up the arithmetic expression) and arithmetic capabilities.
The only dataset on which OccamLlama struggles is GSM8K [40]. We believe the reason for this is primarily because this data involves a large number of fractions and percentages, which Llama does not convert to decimal format. As such, Llama often calls OccamNet with expressions such as “multiplying by 3/4 gives” or “5% of this gives.” Because the switch is not trained on many examples like these in which the number is not in decimal format, it does not realize that OccamNet cannot be used in these cases. Therefore, the switch triggers OccamNet, which is not capable of performing the correct operation because it requires two steps. Future work could address this issue by training the router with more data on this type of situation or by training an OccamLlama with a two layer OccamNet.
Relative error is another important metric that complements accuracy. It measures by how much the answer differs from the true result. For two models with a similar accuracy metric, the relative error they achieve can be very different. Table 6 shows the relative error for the arithmetic experiments. An answer marked correct can have a nonzero relative error because of machine precision limits and because the answer does not report an infinite number of digits.
Interestingly, Llama 2 performs exceptionally poorly on division. By examining outputs, we see that this is because Llama 2 produces an approximately correct output but with the decimal place in the wrong position, leading to a result that is off by many orders of magnitude.
@@
Table 5: Accuracy on reasoning tasks. Higher is Better. Bold indicates best performance.
@@
Average 85.8$\pm$0.5 53.3$\pm$0.7 82.1$\pm$0.6 86.5$\pm$0.5 95.2$\pm$0.3 93.4$\pm$0.7
Table 6: Relative error on arithmetic tasks (%). Lower is Better.
@@
Appendix B Example OccamLlama Generations
[…]
By chance, all six responses happen to be correct.
B.1 MATH401
Example 1
Prompt:
7^4=
Response:
2401. 7 ^4 = 7 × 7 × 7 × 7 = 2401
Example 2
Prompt:
6**3=
Response:
216. 6 ^3 = 6 × 6 × 6 = 216
Example 3
Prompt:
63.0+0.074=
Response:
63.074. 63.0 + 0.074 = 63.074. So the answer is 63.074.
B.2 MultiArith-Float
Example 1
Prompt:
Mike had 30 video games but 6 of them weren’t working. If he wanted to sell the working games for 85.53 each, how much money could he earn?
Response:
Mike had 30 video games. 6 weren’t working, so he had 30 - 6 = 24. He can sell 24 games for 85.53 each. 24 x 85.53 is 2052.720. So Mike could earn 2052.72 dollars.
Example 2
Prompt:
While playing a trivia game, Adam answered 6 questions correct in the first half and 3 questions correct in the second half. If each question was worth 59.08 points, what was his final score?
Response:
Adam got 6 questions correct in the first half, and 3 questions correct in the second half. So he got 6 + 3 = 9. Each question was worth 59.08 points. So his score is 9 * 59.08 = 531.720. The answer is 531.72.
Example 3
Prompt:
Lana’s favorite band was holding a concert where tickets were 181.92 dollars each. Lana bought 5 tickets for herself and her friends and 4 extra tickets in case anyone else wanted to go. How much did she spend?
Response:
Lana bought 5 tickets for herself and her friends. Then 4 extra tickets. So in total she bought 5 + 4 = 9. Each ticket was 181.92 dollars. So 9 tickets will be 9 x 181.92 = 1637.280. The answer is 1637.280.
Appendix C Alternative Architectures and Losses
C.1 Alternative Architectures